OpenStudy (anonymous):

Factoring out the greatest common binomial factor. For question like x(x+6)+5(x+6) the answer would be (x+6) (X+5) but what about questions like x(y-2)+3(2-y)?

Directrix (directrix):

x(y-2)+3(2-y) Factor -1 out of (2 - y) so that it becomes -1(y - 2). Note that -1(y - 2) = (2 - y).

Directrix (directrix):

x(y-2)+3(2-y) = x ( y - 2 ) -3 ( y - 2) = (y - 2) ((x - 3) @7410852963nm